OpenStudy (anonymous):

WILL GIVE MEDAL, FAN, AND TESIMONIAL! Soda Tak claims that Diet Tak has 40mg of sodium per can. You work for a consumer organization that tests such claims. You take a random sample of 60 cans and find that the mean amount of sodium in the sample is 42.4mg. The population standard deviation in all cans is 5.2mg. You suspect that there is more than 40mg of sodium per can. Find the z-score 1.785 2.582 3.575 5.143

OpenStudy (anonymous):

I know that \[z = \chi - \mu / \sigma \] from my AP Statistics notes

OpenStudy (anonymous):

where \[\mu = mean \] and \[\sigma = standard deviation \]

OpenStudy (anonymous):

@ybarrap @agent0smith @amistre64

OpenStudy (anonymous):

@sdfgsdfgs Thank you for looking at my question... For some reason my calculator keeps coming up with error...

OpenStudy (sdfgsdfgs):

sorry my stats is not very good...

OpenStudy (anonymous):

that's ok.

OpenStudy (sdfgsdfgs):

but i found the sol'n 4 u anyway :P it has been asked n answered on OS be4....scroll half way down this post to see the formula plz http://openstudy.com/study#/updates/5339edfde4b0720dc69b6621 in ur case, there IS a right ans ;)

OpenStudy (anonymous):

\[Z = X- E[X] / \sigma (X) / \sqrt{n}\]

OpenStudy (anonymous):

is the equation given.... Sorry I was having dinner with my family we are not allowed to have any electronics at the table... House Rule

OpenStudy (sdfgsdfgs):

hehehee strict family dinning rulez huh? :) its ok but yes, dats the eqn to calculate the Z-score in this case n it works!

OpenStudy (anonymous):

Yup. and X is = score Correct?

OpenStudy (anonymous):

What do X an n stand for?

OpenStudy (sdfgsdfgs):

X=experimental mean and n is sample size

OpenStudy (anonymous):

ok.

OpenStudy (anonymous):

give me a sec to build the equation

OpenStudy (anonymous):

\[Z= 40 - E [42.4] /\sigma [40] / \sqrt{5.2}\]

OpenStudy (anonymous):

I have no idea if I did this right.... I am really bad with setting up my equations

OpenStudy (sdfgsdfgs):

n is not equal to 5.2....

OpenStudy (anonymous):

what are E and \[\sigma \]

OpenStudy (anonymous):

AAAAHHHH!!!! \[\sigma = 5.2\]

OpenStudy (anonymous):

n = 60

OpenStudy (anonymous):

@sdfgsdfgs better

OpenStudy (anonymous):

is E= 40?

OpenStudy (sdfgsdfgs):

:) all correct!

OpenStudy (anonymous):

ok so......

OpenStudy (sdfgsdfgs):

hahahaa cant do this w/o a calculator....u done w dinner yet? ;)

OpenStudy (anonymous):

yup.

OpenStudy (anonymous):

i'm setting it up and plugging it in now

OpenStudy (sdfgsdfgs):

in ur prob, there IS a right ans...unlike the last time...

OpenStudy (anonymous):

\[Z = 40 - 40 (42.4) / (5.2) (40) / \sqrt{60}\]

OpenStudy (anonymous):

is that correct?

OpenStudy (sdfgsdfgs):

Z=X−E[X]/σ(X)/√n is really Z=(X-E)/σ/√n but ur eqn will work out as the 40 above n below will cancel each other out

OpenStudy (anonymous):

ok. so i keep getting either -1644.251219 OR 39.00692735

OpenStudy (sdfgsdfgs):

Z=|(42.4-40)|/(5.2)/√60

OpenStudy (anonymous):

41.40692735

OpenStudy (anonymous):

@kropot72 Can you see what I am doing wrong?

OpenStudy (sdfgsdfgs):

how about this: Z=|(42.4-40)|/(5.2/√60)

OpenStudy (anonymous):

same answer...

OpenStudy (anonymous):

That is the same problem as before

OpenStudy (sdfgsdfgs):

Z=\[\frac{ 42.4-40 }{ \frac{ 5.2 }{ \sqrt{60} } }\]

OpenStudy (anonymous):

3.575

OpenStudy (anonymous):

so the answer would be C Thank you very much for your help!!

OpenStudy (sdfgsdfgs):

welcome n u may want to re-trace ur calculation to see where u made a mistake n did not come up w the right ans. good luck! :)